OpenStudy (anonymous):

Simplify each of the following expressions 8. 47-95-(-73) 10. (2y^6)^3 11. (-3y)(2y^2)

OpenStudy (agent47):

8)47-95+73 = 25 10) 8y^18 11) -6y^3

OpenStudy (anonymous):

can you please support your answer by showing work?

OpenStudy (agent47):

10)2^3*y^(6*3)=8*y^18

OpenStudy (agent47):

11) (-3)*2*y^(1+2)=-6y^3

OpenStudy (mathlegend):

This is also a good site for these rules. Comes with examples and all. http://www.purplemath.com/modules/exponent.htm

OpenStudy (anonymous):

Agent47 what about #8?

OpenStudy (mathlegend):

8. Is just using order of operations addition and subtraction remember?

OpenStudy (mathlegend):

47 - 95 = -48 -48 + 73 = 25

OpenStudy (mathlegend):

We added 73 since two minus signs turn into one big PLUS sign

OpenStudy (mathlegend):

We moved from left to right the whole problem

OpenStudy (mathlegend):

Anymore?

OpenStudy (anonymous):

umm lol yeah actually

OpenStudy (anonymous):

The same rule okay....

OpenStudy (anonymous):

12. (2x^3-9)+(5x^3-10x^2+10) 13. (-6y^2-y+3)-(y^3-10y^2+y+3) 14. (2n+3)(n-6) 15. (x-3)^2

OpenStudy (mathlegend):

So we will have to foil for all of these

OpenStudy (anonymous):

huh? whats foil?

OpenStudy (mathlegend):

Foil explained here: http://www.youtube.com/watch?v=65BAsV0SK3k 2x^3 * 5x^3 = 10x^6 2x^3 * -10x^2 = -20x^5

OpenStudy (mathlegend):

2x^3 * 10 = 20x^3

OpenStudy (mathlegend):

Now we use the -9 -9 * 5x^3= -45x^3 -9 * -10x^2 = 90x^2 -9 * 10 = -90

OpenStudy (mathlegend):

All of the products combined would equal 10x^6-20x^5+20x^3-45x^3+90x^2-90

OpenStudy (mathlegend):

Now we must combine what we can

OpenStudy (agent47):

Good job Legend!

OpenStudy (mathlegend):

So we get 10x^6-20x^5-25x^3+90x2-90

OpenStudy (mathlegend):

Thank You Agent47

OpenStudy (anonymous):

Thanks You Guys! :)

OpenStudy (mathlegend):

13. -6y^2 * y3 = -6y^5 -6y^2*-10y^2= 60y^4 -6y^2 * y = -6y^3 -6y^2 * 3 = -18y^2

OpenStudy (mathlegend):

-y * y^3 = -y^4 -y * -10y^2 = 10y^3 -y * y = -y^2 -y * 3 = -3y

OpenStudy (mathlegend):

3 * y^3 = 3y^3 3 * -10y^2 = 30y^2 3 * y = 3y 3 * 3 = 9

OpenStudy (anonymous):

man this is alot of work Thanks ALOT :)

OpenStudy (mathlegend):

Now we must combine like turns

OpenStudy (mathlegend):

6y^5-61y^4-7y^3+49y^2-9

OpenStudy (mathlegend):

14. We must FOIL 2n * n = 2n^2 2n * -6 = -12n 3 * n = 3n 3 * -6 = -18

OpenStudy (mathlegend):

2n^2 - 12n + 3n - 18

OpenStudy (mathlegend):

Now we combine like terms -12n + 3n = -9n

OpenStudy (mathlegend):

Answer: 2n^2-9n-18

OpenStudy (mathlegend):

Now if I were you I would practice this a lot. In the future your teacher will have you do problems that make you start with: 2n^2-9n-18 and tell you to factor it and get an answer of (2n+3)(n-6)

OpenStudy (mathlegend):

(x-3)^2 notice this is just like saying 7^2

OpenStudy (mathlegend):

7^2 is (7)(7) right?

OpenStudy (mathlegend):

(x-3)(x-3) x*x = x^2 x*-3 = -3x -3*x = -3x -3*-3 = 9

OpenStudy (mathlegend):

Now we can combine like terms

OpenStudy (mathlegend):

x^2-6x+9

OpenStudy (farmdawgnation):

Just a general public service announcement for the crowd. Please don't just give out answers. Rather guide the asker to the answer as per the Code of Conduct. Thanks.

OpenStudy (mathlegend):

ok will do

OpenStudy (mathlegend):

If you need help with anymore I'll guide you through it

OpenStudy (mathlegend):

Just remember problems like these will come up in school again. However they will give you the answer and tell you to work backwards by factoring. :) Factoring Quadratics

OpenStudy (mathlegend):

Make sure you practice a lot.